Consider two normalized eigen functions ψ1 and ψ2 , corresponding to the same eigen value. If




∫ψ∗1ψ2dτ=d ,




where d is real




Find a normalized linear combinations of ψ1 and ψ2 that are orthogonal to




(a) ψ1




(b) ψ1+ψ2




Note: The coefficients of the linear combinations need not be real

1
Expert's answer
2022-03-25T05:47:03-0400

ψ(x) = ψ1(x) + ψ2(x)

ψ+ = 1/1.41 * (ψ1 + xψ2);

ψ- = 1/1.41 * (ψ1 - xψ2);

d/dT = "\int" (ψ1 + ψ2) * -(ψ2)